OpenStudy (anonymous):

If a fly is buzzing randomly around a room 8 ft long, 12 ft wide and 10 ft high, what is the probability that, at any given time, the fly is within 6 feet of the ceiling? Express your answer as a common fraction.

OpenStudy (anonymous):

plz help!

OpenStudy (anonymous):

total area is 8 * 12 * 10 = 960 within 6 feet area is 8 * 12 * 6

OpenStudy (anonymous):

ok sooo

OpenStudy (anonymous):

8 * 12 * 6 = 576 \[\frac{8 * 12 * 6}{8*12*10} = \frac{6}{10} = \frac{3}{5}\]

OpenStudy (anonymous):

so its \[3 \over 5\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

Once again, i will check

OpenStudy (anonymous):

Press good answer if im correct

OpenStudy (anonymous):

I am happy for you :)